Circular No. 6390

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 T. Takeshima, J. K. Cannizzo, and R. Corbet, Goddard Space Flight

Center (GSFC) and Universities Space Research Association; and F. E. Marshall, GSFC, report: “The Rossi X-ray Timing Explorer (RXTE) observed GRS 1739–278 during Mar. 31.754–31.862 UT. The pointing position was slightly offset (to α = 17^{h}42^{m}05^{s}, δ = −27^{o}37′.2, equinox 2000.0) in order to reduce possible contamination from the nearby bright source GRO 1744–28. The source flux in the PCA band (2–60 keV) was 490 mCrab, corrected for the offset from the source, and the spectrum is very soft. There were no detectable oscillations with periods between ∼ 0.002 and ∼ 256 s. In addition, GRS 1739–278 was detected during several slews between other targets, and the following fluxes were found: Mar. 16, 810 mCrab; 22, 430; Apr. 21, 300. Errors on the fluxes from the slew data are conservatively estimated to be 10 percent. After standard processing, data from the pointed observation of GRS 1739–278, as well as other public domain XTE observations, are available over the Internet. Information on how to obtain these

data can be found at http://heasarc.gsfc.nasa.gov/0/docs/xte/xte

1st.html/ (see ‘TOO News’ for a spectrum and lightcurve).”

X-RAY SOURCES A. M. Levine, H. Bradt, D. Chakrabarty, W. Cui, J. G. Jernigan, E.

Morgan, R. Remillard, R. Shirey, and D. A. Smith, Center for Space Research and Department of Physics, Massachusetts Institute of Technology, report that five galactic and Large Magellanic Cloud x-ray sources, in preliminary results from the RXTE All Sky Monitor (ASM) data, appear to have intensified in the band 2–10 keV since late March. The blackhole candidate LMC X-3 has increased from ∼ 15 mCrab in late March to ∼ 40 mCrab. The recurrent transient 4U 1630–47 intensified from < 20 mCrab to ∼ 200 mCrab around Mar. 20 and remains steady at that level. The strength of burster and QPO source 4U 1705–44 has grown linearly with time from Mar. 20 to Apr. 28, at which time it had a flux of ∼ 200 mCrab. The rapid burster MXB 1730–335 suddenly appeared in outburst around Apr. 18, when it appeared to peak at ∼ 200 mCrab; it has weakened since the peak to a level of ∼ 100 mCrab. Finally, the ASM also detects GRS 1739–278 (IAUC 6348) at a level of ∼ 250 mCrab in early Mar., reaching a peak of ∼ 800 mCrab around Mar. 20, and declining to an intensity of ∼ 350 mCrab in late April.

1996 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green